### v2

@@ 6.1 Individual polychoric correlation coefficient
For misspecification fraction $\varepsilon\in\{0,0.1,0.2\}$, we sample $N=1,000$ ordinal responses from this data generating process and estimate the true parameter $\bm{\theta}_{*}$ with the MLE (Olsson,, 1979), our proposed estimator with tuning constant set to $c=0.6$, and, for comparison, the Pearson sample correlation calculated on the integer-valued responses. We repeat this procedure for 1,000 simulated datasets. As performance measures, we calculate the average bias, standard deviation across repetitions, coverage, and length of confidence intervals at significance level $\alpha=0.05$. The coverage is defined as proportion of $(1-\alpha)$-th confidence intervals $[\widehat{\rho}_{N}\mp q_{1-\alpha/2}\cdot\text{SE}(\widehat{\rho}_{N})]$ that contain the true $\rho_{*}$, where $q_{1-\alpha/2}$ is the $(1-\alpha/2)$-th quantile of the standard normal distribution and $\text{SE}(\widehat{\rho}_{N})$ is the standard error of $\widehat{\rho}_{N}$, which is constructed using the limit theory developed in Theorem 1 in Appendix A. The length of a confidence interval is given by $2\cdot q_{1-\alpha/2}\cdot\text{SE}(\widehat{\rho}_{N})$.
@@
$\varepsilon=0$ Polycor robust 0.504 0.004 0.027 0.930 0.104
Polycor ML 0.500 0.000 0.026 0.943 0.102
Sample cor 0.457 $-0.043$ 0.025 0.702 0.110
$\varepsilon=0.1$ Polycor robust 0.466 $-0.034$ 0.038 0.911 0.152
Polycor ML 0.097 $-0.403$ 0.029 0.000 0.134
Sample cor 0.084 $-0.416$ 0.026 0.000 0.124
$\varepsilon=0.2$ Polycor robust 0.439 $-0.061$ 0.051 0.951 0.220
Polycor ML $-0.172$ $-0.672$ 0.028 0.000 0.133
Sample cor $-0.133$ $-0.633$ 0.026 0.000 0.123
@@
Figure 4 visualizes the bias of each estimator with respect to the true polychoric correlation $\rho_{*}$ across the  1,000 simulated datasets. An analogous plot for the whole parameter $\bm{\theta}_{*}$ can be found in Appendix C; the results are similar to those of $\rho_{*}$. For all considered misspecification fractions, the estimates of the MLE and sample correlation are similar, which is expected because these two estimators are known to yield similar results when there are five or more rating options and the discretization thresholds are symmetric (cf., Rhemtulla et al.,, 2012). In the absence of misspecification, both MLE and the robust estimator yield accurate estimates. Both estimates are nearly equivalent to one another in the sense that their point estimates, standard deviation, and coverage at significance level $\alpha=0.05$ are very similar (Table 1). However, when misspecification is introduced, MLE, sample correlation, and robust estimator yield noticeably different results. At misspecification degree $\varepsilon=0.1$, MLE and sample correlation are substantially biased with average estimates of $0.097$ and $0.084$, corresponding to biases of $-0.403$ and $-0.416$, respectively, as well as zero coverage. In contrast, the robust estimator maintains accuracy with an average estimate of $0.466$, which corresponds to only a minor bias of $-0.034$ and a good coverage of $0.911$ (Table 1). When the misspecification is increased to $\varepsilon=0.2$, the contrast between the estimators becomes even stronger. While the robust estimator is still remarkably close to the truth with a small bias of $-0.061$, MLE and sample correlation produce estimates that are not only severely biased (biases of $-0.672$ and $-0.633$), but also sign-flipped: While the true correlation is strongly positive (0.5), both estimates are considerably negative ($-0.172$ and $-0.133$). It is worth noting that in the presence of misspecification, the confidence intervals of the robust estimator are wider than those of the MLE (see Table 1). This is expected because of a well-known trade-off between robustness and efficiency: An estimator that is designed to reduce bias, like a robust estimator, will inevitably have a larger estimation variance (e.g., Huber & Ronchetti,, 2009). These wider confidence intervals furthermore explain why the robust estimator improves its coverage in Table 1 when the misspecification fraction is increased from 0.1 to 0.2.
@@ 6.2 Polychoric correlation matrix
Following the five-dimensional simulation design in Foldnes & Grønneberg, (2020), there are $q=5$ ordinal variables with polychoric correlation matrix as in Table 5 such that the pairwise correlations vary from a low 0.2 to a moderate 0.56.333The correlation matrix in Table 2 has the additional interpretation of being the covariance matrix of a factor model for a single factor with loadings vector $(0.8,0.7,0.6,0.5,0.4)^{\top}$. For all latent variables, the discretization thresholds are set to, in ascending order, $\Phi_{1}^{-1}(0.1)=-1.28,\ \Phi_{1}^{-1}(0.3)=-0.52,\ \Phi_{1}^{-1}(0.7)=0.56$, and $\Phi_{1}^{-1}(0.9)=1.28$, such that each ordinal variable can take five possible values. A visualization of the implied distribution of each ordinal variable can be found in Figure 5 in Foldnes & Grønneberg, (2020).
@@
For misspecification fraction $\varepsilon\in\{0,0.1,0.2\}$, we sample $N=1,000$ ordinal five-dimensional responses from this data generating process and use them to estimate the polychoric correlation matrix in Table 2 via ML and our robust estimator (again with tuning constant $c=0.6$). We repeat this procedure 1,000 times.
@@ 7.1 Background and study design
Arias et al., (2020) collect measurements of three Big Five traits in this way, namely extroversion, neuroticism, and conscientiousness. [^4] Arias et al., (2020) synonymously refer to neuroticism as emotional stability. Furthermore, in addition to the three listed traits, Arias et al., (2020) collect measurements of the trait dispositional optimism by using a different instrument, and another scale that is designed to not measure any construct. We do not consider these scales in this empirical demonstration. Furthermore, the data of Arias et al., (2020) has been made publicly available by the authors at https://osf.io/n6krb. The sample that we shall use, Sample 1 in Arias et al., (2020), consists of $N=725$ online respondents who are all U.S. citizens, native English speakers, and tend to have relatively high levels of reported education (about 90% report to hold an undergraduate or higher degree). Concerned about respondent inattention in their data, Arias et al., (2020) construct a factor mixture model for detecting inattentive/careless participants. Their model crucially relies on response inconsistencies to polar opposite adjectives and is designed to primarily detect careless straightlining responding. They find that careless responding is a sizable problem in their data. Their model finds evidence of straightliners, and the authors conclude that if unaccounted for, they can substantially deteriorate the fit of theoretical models, produce spurious variance, and overall jeopardize the validity of research results.
@@ 8.1 Preliminaries
The simulation studies in Section 6 have been concerned with a situation in which the polychoric model is misspecified for a subset of a sample (partial misspecification). However, as outlined in Section 4.3, another misspecification framework of interest is that of distributional misspecification where the model is misspecified for the entire sample. Suppose that instead of a bivariate standard normal distribution, the latent variables $(\xi,\eta)$ jointly follow an unknown and unspecified distribution $G$. In this framework, the object of interest is the population correlation between $\xi$ and $\eta$ under distribution $G$, that is, $\rho_{G}=\mathrm{Cor}_{G}\left[\xi,\ \eta\right]$, rather than a polychoric correlation coefficient.
Although distributional misspecification is not covered by the misspecified distribution (7) under which we study the theoretical properties of our proposed estimator, our estimator could still offer a gain in robustness compared to ML estimation under distributional misspecification. Specifically, the robust estimator may still be useful if the central part of the nonnormal distribution $G$ is not too different from a standard bivariate normal distribution: Intuitively, if the difference between $G$ and a standard normal distribution is mainly in the tails, $G$ can be approximated by a mixture of a standard normal distribution (to cover the central part) and some other distribution $H$ (to cover the tails). The ML estimator tries to treat influential observations from the tails—which cannot be fitted well by the polychoric model—as if they were normally distributed, resulting in a possibly large estimation bias. In contrast, the robust estimator uses the normal distribution only for observations from the central part—which may fit the polychoric model well enough—and downweights observations from the tails. Thus, as long as such a mixture is a decent approximation of the nonnormal distribution $G$, the robust estimator should perform reasonably well. However, if $G$ cannot be approximated by such a mixture, neither the ML estimator nor our estimator can be expected to perform well. Overall, though, our estimator could offer an improvement in terms of robustness to distributional misspecification.

### v3

@@ 6.1 Individual polychoric correlation coefficient
For misspecification fraction $\varepsilon\in\{0,0.01,0.05,0.1,0.15,0.2\}$, we sample $N=1,000$ ordinal responses from this data generating process and estimate the true parameter $\bm{\theta}_{*}$ with the MLE (Olsson,, 1979), our proposed estimator with tuning constant set to $c=0.6$, and, for comparison, the Pearson sample correlation calculated on the integer-valued responses. We repeat this procedure for 1,000 simulated datasets. As performance measures, we calculate the average bias, standard deviation across repetitions, coverage, and length of confidence intervals at significance level $\alpha=0.05$. The coverage is defined as proportion of $(1-\alpha)$-th confidence intervals $[\widehat{\rho}_{N}\mp q_{1-\alpha/2}\cdot\text{SE}(\widehat{\rho}_{N})]$ that contain the true $\rho_{*}$, where $q_{1-\alpha/2}$ is the $(1-\alpha/2)$-th quantile of the standard normal distribution and $\text{SE}(\widehat{\rho}_{N})$ is the standard error of $\widehat{\rho}_{N}$, which is constructed using the limit theory developed in Theorem 1 in Appendix A. The length of a confidence interval is given by $2\cdot q_{1-\alpha/2}\cdot\text{SE}(\widehat{\rho}_{N})$.
@@
$\varepsilon=0$ Polycor robust 0.505 0.005 0.027 0.942 0.104
Polycor ML 0.501 0.001 0.027 0.949 0.102
Sample cor 0.457 $-0.043$ 0.025 0.686 0.110
$\varepsilon=0.01$ Polycor robust 0.497 $-0.003$ 0.029 0.944 0.110
Polycor ML 0.447 $-0.053$ 0.027 0.528 0.109
Sample cor 0.408 $-0.092$ 0.025 0.073 0.113
$\varepsilon=0.05$ Polycor robust 0.479 $-0.021$ 0.033 0.912 0.130
Polycor ML 0.272 $-0.228$ 0.029 0.000 0.126
Sample cor 0.242 $-0.258$ 0.026 0.000 0.120
$\varepsilon=0.1$ Polycor robust 0.466 $-0.034$ 0.039 0.906 0.153
Polycor ML 0.096 $-0.404$ 0.029 0.000 0.134
Sample cor 0.083 $-0.417$ 0.026 0.000 0.124
$\varepsilon=0.15$ Polycor robust 0.456 $-0.044$ 0.044 0.922 0.176
Polycor ML $-0.050$ $-0.550$ 0.029 0.000 0.136
Sample cor $-0.038$ $-0.538$ 0.02y 0.000 0.124
$\varepsilon=0.2$ Polycor robust 0.440 $-0.060$ 0.052 0.943 0.220
Polycor ML $-0.173$ $-0.673$ 0.028 0.000 0.133
Sample cor $-0.134$ $-0.634$ 0.026 0.000 0.123
@@
Figure 4 visualizes the bias of each estimator with respect to the true polychoric correlation $\rho_{*}$ across the  1,000 simulated datasets. An analogous plot for the whole parameter $\bm{\theta}_{*}$ can be found in Appendix C; the results are similar to those of $\rho_{*}$. For all considered misspecification fractions, the estimates of the MLE and sample correlation are similar, which is expected because these two estimators are known to yield similar results when there are five or more rating options and the discretization thresholds are symmetric (cf., Rhemtulla et al.,, 2012). In the absence of misspecification, both MLE and the robust estimator yield accurate estimates. Both estimates are nearly equivalent to one another in the sense that their point estimates, standard deviation, and coverage at significance level $\alpha=0.05$ are very similar (Table 1). However, when misspecification is introduced, MLE, sample correlation, and robust estimator yield noticeably different results. Already at the small misspecification fraction $\varepsilon=0.01$ (corresponding to only 10 observations), MLE and sample correlation are noticeably biased, resulting in poor coverage of only about 0.53 and 0.07, respectively. Increasing the misspecification fraction to the still relatively small value of $\varepsilon=0.05$, MLE and sample correlation start to be substantially biased with average biases of $-0.23$ and $-0.26$, respectively, leading to zero coverage. The biases of these two methods further deteriorates as the misspecification fraction is gradually increased. At $\varepsilon=0.2$, MLE and sample correlation produce estimates that are not only severely biased (biases of $-0.67$ and $-0.63$), but also sign-flipped: While the true correlation is strongly positive (0.5), both estimates are considerably negative ($-0.17$ and $-0.13$). In stark contrast, the proposed robust estimator remains accurate throughout all considered misspecification fractions: At the small $\varepsilon=0.01$, the robust estimator is nearly unaffected, while at the larger $\varepsilon=0.2$, it only exhibits a minor bias of about $-0.06$. In all considered scenarios, its coverage remained above 0.9. It is worth noting that in the presence of misspecification, the confidence intervals of the robust estimator are wider than those of the MLE (see Table 1). This is expected because of a well-known trade-off between robustness and efficiency: An estimator that is designed to reduce bias, like a robust estimator, will inevitably have a larger estimation variance (e.g., Huber & Ronchetti,, 2009). These wider confidence intervals furthermore explain why the robust estimator improves its coverage in Table 1 when the misspecification fraction is increased from 0.1 to 0.15 and 0.2.
@@ 6.2 Polychoric correlation matrix
Following the five-dimensional simulation design in Foldnes & Grønneberg, (2020), there are $q=5$ ordinal variables with polychoric correlation matrix as in Table 5 such that the pairwise correlations vary from a low 0.2 to a moderate 0.56.444The correlation matrix in Table 2 has the additional interpretation of being the covariance matrix of a factor model for a single factor with loadings vector $(0.8,0.7,0.6,0.5,0.4)^{\top}$. For all latent variables, the discretization thresholds are set to, in ascending order, $\Phi_{1}^{-1}(0.1)=-1.28,\ \Phi_{1}^{-1}(0.3)=-0.52,\ \Phi_{1}^{-1}(0.7)=0.56$, and $\Phi_{1}^{-1}(0.9)=1.28$, such that each ordinal variable can take five possible values. A visualization of the implied distribution of each ordinal variable can be found in Figure 5 in Foldnes & Grønneberg, (2020).
@@
For misspecification fraction $\varepsilon\in\{0,0.01,0.05,0.1,0.15,0.2\}$, we sample $N=1,000$ ordinal five-dimensional responses from this data generating process and use them to estimate the polychoric correlation matrix in Table 2 via ML and our robust estimator (again with tuning constant $c=0.6$). We repeat this procedure 1,000 times.
@@ 7.1 Background and study design
Arias et al., (2020) collect measurements of three Big Five traits in this way, namely extroversion, neuroticism, and conscientiousness. [^5] Arias et al., (2020) synonymously refer to neuroticism as emotional stability. Furthermore, in addition to the three listed traits, Arias et al., (2020) collect measurements of the trait dispositional optimism by using a different instrument, and another scale that is designed to not measure any construct. We do not consider these scales in this empirical demonstration. Furthermore, the data of Arias et al., (2020) have been made publicly available by the authors at https://osf.io/n6krb. The sample that we shall use, Sample 1 in Arias et al., (2020), consists of $N=725$ online respondents who are all U.S. citizens, native English speakers, and tend to have relatively high levels of reported education (about 90% report to hold an undergraduate or higher degree). Concerned about respondent inattention in their data, Arias et al., (2020) construct a factor mixture model for detecting inattentive/careless participants. Their model crucially relies on response inconsistencies to polar opposite adjectives and is designed to primarily detect careless straightlining responding. They find that careless responding is a sizable problem in their data. Their model finds evidence of straightliners, and the authors conclude that if unaccounted for, they can substantially deteriorate the fit of theoretical models, produce spurious variance, and overall jeopardize the validity of research results.
@@ 8.1 Preliminaries
The simulation studies in Section 6 have been concerned with a situation in which the polychoric model is misspecified for a subset of a sample (partial misspecification). However, as outlined in Section 4.3, another misspecification framework of interest is that of distributional misspecification where the model is misspecified for the entire sample. Suppose that instead of a bivariate standard normal distribution, the latent variables $(\xi,\eta)$ jointly follow an unknown and unspecified distribution $G$. In this framework, the object of interest is the population correlation between $\xi$ and $\eta$ under distribution $G$, that is, $\rho_{G}=\mathrm{Cor}_{G}\left[\xi,\ \eta\right]$, rather than a polychoric correlation coefficient. Estimators for such situations where the population distribution $G$ is nonnormal have been proposed by Lyhagen & Ornstein, (2023) and Jin & Yang-Wallentin, (2017).
Huber & Ronchetti, (2009, p. 4) note that, although conceptually distinct, robustness to distributional misspecification and partial misspecification are “practically synonymous notions”. Hence, despite distributional misspecification not being covered by the misspecified distribution (7) under which we study the theoretical properties of our proposed estimator, our estimator could still offer a gain in robustness compared to ML estimation under distributional misspecification. Specifically, the robust estimator may still be useful if the central part of the nonnormal distribution $G$ is not too different from a standard bivariate normal distribution: Intuitively, if the difference between $G$ and a standard normal distribution is mainly in the tails, $G$ can be approximated by a mixture of a standard normal distribution (to cover the central part) and some other distribution $H$ (to cover the tails). The ML estimator tries to treat influential observations from the tails—which cannot be fitted well by the polychoric model—as if they were normally distributed, resulting in a possibly large estimation bias. In contrast, the robust estimator uses the normal distribution only for observations from the central part—which may fit the polychoric model well enough—and downweights observations from the tails. Thus, as long as such a mixture is a decent approximation of the nonnormal distribution $G$, the robust estimator should perform reasonably well. However, if $G$ cannot be approximated by such a mixture, neither the ML estimator nor our estimator can be expected to perform well. Overall, though, our estimator could offer an improvement in terms of robustness to distributional misspecification.